Subject: 3.5-m Users Committee minutes 3/19/07

From: Bruce Gillespie

Submitted: Mon, 19 Mar 2007 15:03:02 -0600

Message number: 1020 (previous: 1019, next: 1021 up: Index)

                         APO 3.5-m Users Committee Phonecon, 3/19/07

Attending:  Suzanne Hawley, Bruce Gillespie, Scott Anderson, John  
Bally, John Fulbright, Al Harper, Remy Indebetouw

Absent: Jon Holtzman, Michael Strauss

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton - (Strauss) [e-mailed from Michael: "For the record, our  
recent observing has been problem-free; nothing
to report on that front from Princeton]

JHU (Fulbright) - John noted that recent JHU programs have had better  
observing weather, and that some student site orientation has occurred.

Chicago (Harper) - Al mentioned that an increase in the Chicago  
demand for the 3.5-m can be attributed to a couple of new observers  
being added at Chicago in the past year, and some new SDSS-inspired  
follow-up observing programs.

Washington (Anderson) - Scott reiterated the desire of their new  
users for a TUI interface to SPIcam.

NMSU - (Holtzman) no report

Virginia (Indebetouw) - Remy said he's received no recent user feedback.

Colorado (Bally) - John talked about the recent fast sub-frame  
observing mode for NIC-FPS, which was implemented to support the  
Pluto occultation.  One glitch discovered is that if you ask for full- 
frame exposures with Fowler sampling set to a large number, the  
instrument control computer crashes.  John is also working on the NIC- 
FPS users manual with Jon Holtzman.


Discussion of telescope/instruments report - Gillespie:

The report is given below; there was only a little discussion.


      3.5-m Telescope, Instruments, and CIF Projects Highlights,  
2/20/07 through 3/16/07

                                  Bruce Gillespie

0) Overview

It has only been three weeks since the last meeting, so there is not  
much to report.  The weather at the site is beginning to be spring- 
like, with some unusually warm temperatures and some windy  
conditions.  The telescope and instruments have been largely well- 
behaved, and the commissioning run of Agile (the UW high-speed  
photometer) has just begun.  Various engineering activities and tests  
have made progress.

1) Telescope

Other than routine maintenance and repairs, not much happened with  
the telescope during the period, and it performed very well.  There  
was an isolated problem with one of the collimation/focus actuators  
on the secondary mirror, and tightening a coupling seemed to fix the  
problem.  Some testing was done related to CIF projects (see below)

2) Instruments

DIS:  The recent troubles with warm-ups on the red camera appear to  
have been solved by the cryotiger work done last month.

NIC-FPS:  Tests were run to enable fast sub-framing operation of the  
detector, in support of the upcoming Pluto occultation observing  

Agile: [in commissioning--report to be given at the meeting by Suzanne]

3) Engineering and CIF projects

Nick McDonald made additional measurements of the repeatability of  
the tertiary mirror rotation as part of the design studies for  
upgrading the tertiary mounting system--it now appears hopeful that  
the tertiary rotation is not making any significant contribution to  
pointing errors.  He also has completed assembling the rebuilt "old"  
telescope drive boxes that are intended as hot-swappable spares for  
the telescope.

Tests of the confocalness of the NA2 guider and the main image focal  
plan were continued, indicating that we may indeed be able to use the  
guider focus to slave the focus of the telescope, aka "auto-focus".   
Some software work needs to be done to continue this project.

A new version of the TCC and TUI were released.

4) Miscellaneous

The summer shutdown plans are firming up, which is being targeted for  
20 August through 16 September (plus three days in July); the work  
list is given in the minutes of the last Users Committee meeting.

The old tower anemometer was repaired (a broken optical fiber), and a  
new "backup" anemometer was installed on the roof of the Baldwin  
Operations Building.  The two devices give different results due to  
their locations (the tower readings are higher), so a calibration  
correction is being researched to apply to the data from the backup  


New plans for NIC-FPS - Bally:

John reported that Stephane Beland is now working on NIC-FPS two days/ 
week--telescope staff and users should send problem reports to both  
John ( and Stephane (

   o The NIR Fabry-Perot etalon and controller have returned from  
England and are back at Boulder, and a dewar has been ordered to test  
the etalon at cryo temperatures.  The etalon controller has been  
modified to have a slower x10 response time, in order to eliminate  
the oscillations seen earlier.  A CU student is lined up to help with  
the tests, which will use the engineering-grade chip with the  
Rockwell controller (the one we didn't choose to use at APO).  If the  
cryo test don't work, they will modify the F-P so that it runs at a  
warmer temperature, 180 to 200 degrees K.

   o The science-grade chip will most likely be sent to Rockwell to  
connect some leads that will enable reading all 16 channels of the  
array.  This will reduce the read noise by allowing much faster  
multiple-read Fowler sampling, which we can't do now with the single- 
channel readout.  Also, the video cards in the Leach controller will  
be replaced to handle the extra channels.  There is some risk in  
having Rockwell work on the chip, but it is considered to be minimal,  
and the engineering-grade chip is a suitable backup if something bad  
happens to the science-grade chip.

   o CU proposes that ARC buy a second Leach controller for NIC-FPS,  
which then can be used for the tests in Boulder this summer, and then  
can be used as a spare at the site for both NIC-FPS and TripleSpec.   
Suzanne said that this proposal will be discussed in April during the  
engineering reviews being held at APO.  John stated that the  
engineering work planned for NIC-FPS during the summer might keep the  
instrument in Boulder until ~October, and he will keep us posted on  
the schedule.


Agile update - Hawley:

Agile, a high-speed blue/visible photometer is at APO for its initial  
commissioning runs.  Everything fit together and it saw first light,  
but a cooling problem with the camera was encountered.  A replacement  
camera is on its way and will hopefully be installed on the telescope  
for tonight's observing.  Anjum Mukadum, Ed Mannery, and Russell  
Owen, all at UWash, are the team putting this instrument and software  
together.  There were no non-Washington requests to use Agile in the  
2nd quarter, but we hope that its usability will be proven to the  
level that it can be requested for use in the 3rd quarter by the  
general user community.

Suzanne said that for a U=12 magnitude star, you can get 20,000  
counts in U in about a second, and the read noise is expected to be  
~10e.  The chip is optimized for the UV and blue, but is also usably  
sensitive in the red.  Anjum will send us a characterization of the  
instrument following its commissioning, which continues this week and  
later in May.


Actions from earlier meetings - group

[open from previous months]:

ACTION:  Rene to collect history of 3.5-m publications, and Bruce  
will make a space for them on the APO website.
STATUS:  Gretchen and Don York have completed a draft the 3.5-m  
publications list.  Once reviewed, it will be added as an appendix to  
the Futures Committee report (already available on the APO web site),  
and the list will also be placed in a stand-alone web location.   
Members of the Users Committee will be invited to check the list  
against known publications from the respective institutions.

[new actions from last month]:

ACTION: It would be good to do some tests to determine the accuracy  
of the [DIS grating] moves so users could evaluate if not checking  
after the move in special situations would be acceptable.
STATUS: Mark Klaene is looking into this and will report back.

ACTION: Suzanne will follow up with Russell Owen about TUI  
scripting.  Users are encouraged to provide feedback to their users  
reps about integrating Spicam into TUI.
STATUS: Suzanne talked to Russell and he is willing to put some  
modest number of scripts into TUI.  He will talk to Jon Holtzman to  
discuss what scripts need to be ported, and how difficult it will  
be.  Other users who use SPIcam scripts are encouraged to contact  
their users committee representative and Russell.


Update on Telescope Systems Workshop - Hawley:

Suzanne added to the previous reports on the Workshop by saying that  
the mid-sized telescope Directors are planning to have a discussion  
about the possibility of future NSF funding in exchange for offering  
community access,
particulary to "unique" capabilities, e.g., NIC-FPS Fabry-Perot.  We  
are well positioned to take advantage of such a program because of  
our well-tested remote observing capability.


Next meeting:

The next Users Committee phonecon will be on Monday, 16 April, at  
8:30 AM Pacific Time.  Agenda and other materials will be sent to the  
committee members during the preceding week.  Topics will include  
reports on TripleSpec, the other engineering reviews being held in  
early April, and planned telescope improvement projects.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1020 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to